OpenStudy (karolynxx):

What is the value of x in the equation 1/5x – 2/3y = 30, when y = 15? 4 8 80 200

OpenStudy (karolynxx):

im sorry had to rewrite the problem forgot the fraction

OpenStudy (lochana):

okay. so the first thing is to make this equation for x. \[\frac{1}{5}x = 30 + \frac{2}{3}y\]

OpenStudy (lochana):

now you can multiply both sides by 5 . \[x = 30*5 + 5*\frac{2}{3}y\]

OpenStudy (lochana):

okay?

OpenStudy (karolynxx):

soo it would be 200?

OpenStudy (lochana):

yes. you are correct